OpenStudy (anonymous):

if you weigh 490N on earth how much do you weigh on the moon?

OpenStudy (anonymous):

\[F=m \times g\] 490N= m * 9'8 (on earth) So m= 50 kg, On the moon the gravity is 1,622m/s^2 F=50 * 1,622 = 81,1N

OpenStudy (michele_laino):

please note that gravity on the moon is g/6, where g=9.81 m/sec^2, so?